OpenStudy (anonymous):

The average of two numbers a and b is given by A = a +b/2 If the average of two numbers is 75, and if one of the numbers is 91, the what is the other number?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {(91 + x) \over 2} = 75 }\) \(\Large \color{purple}{\rightarrow 91 + x = 150 }\) \(\Large \color{purple}{\rightarrow x = 150 - 91 = 59 }\)